OpenStudy (anonymous):

6. If the lateral area of a right circular cylinder is 9m^2 and its volume is 27m^3, find its diameter.

OpenStudy (anonymous):

Can you answer this problem?

OpenStudy (michele_laino):

Plese note that if I call A as the lateral area of your right circular cylinder, and V its volume, I can write: \[A=2 \pi *r*h\] and: \[V=\pi*r ^{2}*h\] where h is the height of your cylinder and r is its radius. Now substituting your numerica data in the above equations, I get: \[2 \pi*r*h=9\] and: \[\pi*r ^{2}*h=27\] So you have to solve those pair of equations, namely a system of two equations in two variable. DO you know how to solve it, please?

OpenStudy (anonymous):

I dont know how..

OpenStudy (anonymous):

Its not clear to me yet.

OpenStudy (anonymous):

Oh! I got the answer already

OpenStudy (michele_laino):

Ok! I help you to solve it. Please solve first equation in order to find h, namely you have to divide both sides of the first equation by 2*pi*r, please try now

OpenStudy (anonymous):

I now got the answer. The radius is 6 and the diameter is 12.

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thank you!

OpenStudy (michele_laino):

thank you!

OpenStudy (anonymous):

A right circular cone is inscribed in a cube whose diagonal measures 8sqrtof3cm. Find the volume of the cone.

OpenStudy (anonymous):

Can u answer this please?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

If I call l the edge of your cube, and d its diagonal, then applyng twice the Pitagora theorem, I get: \[d=l*\sqrt{3}\] or squaring both sides: \[d ^{2}=3*l ^{2}\] Now substituting your numerical dat, I find: \[l=2\sqrt{2}\] Radius R of your cone is equal to l/2, so: \[R=\sqrt{2}\] height H of your cone is: \[H=l=2\sqrt{2}\] Finally applying the formula to find the volume of a cone, we can write: \[V=\frac{ 1 }{ 3 }\pi*R^{2}*H=\frac{ \pi }{ 3 }*2*2\sqrt{2}=\frac{ 4 \pi }{ 3 }\sqrt{2}\]

OpenStudy (michele_laino):

@Paulaaquino please, check all of my statements!

OpenStudy (michele_laino):



OpenStudy (anonymous):

I got the volume already. and its 134.04 but i used another solution (;

OpenStudy (michele_laino):

you are right, I have made an error, sorry, I write all corrections. whit same meaning of symbols, I write: \[l=8, R=4V=\frac{ 1 }{ 3 }*\pi 16*8\] check this, I think it is correct!

OpenStudy (michele_laino):

@Paulaaquino is it correct?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Its okay, we are all persons we make mistakes :)

OpenStudy (anonymous):

9. The base of the right prism is a regular pentagon with one side measuring 12cm. find the volume and the total surface area of the altitude measures 48 cm.

OpenStudy (michele_laino):

thank you!

OpenStudy (anonymous):

Can you try this one?

OpenStudy (michele_laino):

Ok!

OpenStudy (anonymous):

I have so many questions because i have an upcoming quiz on tuesday so i need to review sorry :)

OpenStudy (michele_laino):

the apothem a, of your pentgon is: a=8.26cm, so the area of your pentagon is: (perimeter*apothema)/2=\[A=\frac{ 12*5*8.26 }{ 2 }=247.8 cm ^{2} \] the lateralarea S is:perimeter*altitude H, so: \[S=12*5*48=2880cm ^{2}\] then the total area S_tis: S+2*A, then: \[S _{t}=S+2*A=2880+2*495.6=3375.6cm ^{2}\] and the volume V, is A*, namely: \[V=A*H=247.8*48=11,894.4cm ^{2}\] Please, check all of my statements above!

OpenStudy (anonymous):

What wll be the right section of then prism? Because as the formula states: LSA= Pe; where P is the perimeter of the right section and e is the length of the lateral edge. I am so confused with the right section term.

OpenStudy (anonymous):

So, meaning to say, the height is equal to the lateral edge?

OpenStudy (anonymous):

Am i correct?

OpenStudy (michele_laino):

I think the righ cross section is a pentagon like the bases. Area of a pentagon is equal, as I said befor, to the product between the perimeter of the pentagon and its apothema

OpenStudy (michele_laino):

I don't understand, what do you means by the "lateral edge", please?

OpenStudy (michele_laino):

Sorry, ...what do you mean...

OpenStudy (anonymous):

The side of the prism

OpenStudy (anonymous):

Lateral, only the side of the prism without the two bases

OpenStudy (michele_laino):

then Ok! the height is the lateral edge

OpenStudy (anonymous):

Oh, so the height and the side of the prism are equal?

OpenStudy (michele_laino):

Not precisely, the height is equal to 48 cm, whereas the side is what I called the edge of each of the bases of your prism and it's equal to 12 cm

OpenStudy (anonymous):

The sides of the base are not the same as the edges of the lateral surface, i guess?